OpenStudy (anonymous):

Consider the function y(x)=2+square root x-1 a) state the domain b) state the range c) draw a reasonable graph of the function

OpenStudy (colouringme):

Okay, I'll try to help you, let's hope I remember this right... First, you need to determine the domain, by eliminating impossible values. you can square root almost all numbers except the ones that have...

OpenStudy (colouringme):

Actually, can you write down the equation again, using the equation button so I would know whether it is \[\sqrt{x-1}\] or \[\sqrt{x} - 1\]

OpenStudy (anonymous):

\[y(x)=2+\sqrt{x-1}\]

OpenStudy (colouringme):

thanks! do you know what is not allowed inside a square root, what value that is unacceptable?

OpenStudy (anonymous):

no it just states consider the function and then the equation no other info is given

OpenStudy (colouringme):

yes, but generally, you cannot obtain the result of square root if the value inside the square root is....

OpenStudy (anonymous):

to be completely honest i have no idea what it is asking

OpenStudy (colouringme):

what am I asking or what the question is about?

OpenStudy (anonymous):

its asking for the domain which has to stay real so all the square root numbers must be real numbers and obviously the range will just come after finding the domain and then i have to graph it based on the domain and range

OpenStudy (colouringme):

yes, so you know what value of x must be, right? what is the value of x, so that the value inside of the square root is no less than 0 ?

OpenStudy (anonymous):

yes because you cant do the square root of a negative number so it all has to be bigger than 0

OpenStudy (colouringme):

so x values are...

OpenStudy (colouringme):

write down it using any of these : > < or =

OpenStudy (anonymous):

wouldnt it be \[x >0\] for the domain

OpenStudy (colouringme):

yes

OpenStudy (anonymous):

wait or would it be \[x \ge1\] because it has to at least be 1 due to the -1

OpenStudy (colouringme):

yes that can do too

OpenStudy (colouringme):

I'd rather use the second one though,

OpenStudy (anonymous):

agreed alright so how to i find the range

OpenStudy (anonymous):

do*

OpenStudy (colouringme):

input the lowest possible value for the x, and you'll find the result, it would be the minimum value... You can also find it using graph (which is sadly, sorry, I quite can't help you with that one)

OpenStudy (anonymous):

alright thank you very much

OpenStudy (colouringme):

umm, have you found the range?

OpenStudy (anonymous):

im finding but my problem was understanding what was being asked so i got it

OpenStudy (colouringme):

ooh what is the answer ( I just want to make sure we're on the same page ;) )

OpenStudy (anonymous):

would it be \[y \ge2\]

OpenStudy (colouringme):

yas it is! (:

OpenStudy (anonymous):

thank you for all your help i really appreciate it

OpenStudy (colouringme):

you're nice and patient too!